# Toward Self-Improvement of LLMs via Imagination, Searching, and Criticizing

### v1

@@ Abstract
Despite the impressive capabilities of Large Language Models (LLMs) on various tasks, they still struggle with scenarios that involves complex reasoning and planning. Recent work proposed advanced prompting techniques and the necessity of fine-tuning with high-quality data to augment LLMs’ reasoning abilities. However, these approaches are inherently constrained by data availability and quality. In light of this, self-correction and self-learning emerge as viable solutions, employing strategies that allow LLMs to refine their outputs and learn from self-assessed rewards. Yet, the efficacy of LLMs in self-refining its response, particularly in complex reasoning and planning task, remains dubious. In this paper, we introduce AlphaLLM for the self-improvements of LLMs, which integrates Monte Carlo Tree Search (MCTS) with LLMs to establish a self-improving loop, thereby enhancing the capabilities of LLMs without additional annotations. Drawing inspiration from the success of AlphaGo, AlphaLLM addresses the unique challenges of combining MCTS with LLM for self-improvement, including data scarcity, the vastness search spaces of language tasks, and the subjective nature of feedback in language tasks. AlphaLLM is comprised of prompt synthesis component, an efficient MCTS approach tailored for language tasks, and a trio of critic models for precise feedback. Our experimental results in mathematical reasoning tasks demonstrate that AlphaLLM significantly enhances the performance of LLMs without additional annotations, showing the potential for self-improvement in LLMs.
@@ 1 Introduction
LLMs, trained on trillions of tokens with billions of parameters have shown unparalleled capabilities in a wide range of natural language processing tasks (Touvron et al., 2023b; Team et al., 2023; OpenAI, 2023). Nevertheless, they continue to face challenges in scenarios requiring complex reasoning and strategic planning  (Valmeekam et al., 2022; Stechly et al., 2024). While advanced prompting approaches such as Chain, Tree, Graph-of-Thought (Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Ding et al., 2023), which generate intermediate steps in the reasoning process demonstrate large improvements on reasoning capability of LLMs, it remains essential to fine-tune LLMs using a substantial volume of high-quality, supervised data to fundamentally improve the model performance (Nye et al., 2021; Lewkowycz et al., 2022; Chung et al., 2022). This methodology is inherently limited by the scope and quality of data that humans can provide.
@@
Considering existing challenges, the concept of self-correction and self-learning have been proposed as promising solutions (Madaan et al., 2024; Saunders et al., 2022; Chen et al., 2024). Within these framework, LLMs typically operate by employing two main strategies: 1) they continuously refine their responses based on the feedback of their past responses, and 2) they extensively sample responses then learn from preferences judged by itself as reward models with PPO or DPO (Yuan et al., 2024a, b; Chen et al., 2024). However, it remains a matter of ongoing research whether LLMs can effectively critique their own outputs to either enhance response quality or apply a scalar reward to indicate the quality of responses, especially in contexts demanding intricate planning and reasoning (Valmeekam et al., 2022; Stechly et al., 2024; Huang et al., 2023; Hong et al., 2023). On the other hand, advanced search algorithms such as Monte Carlo Tree Search (MCTS), combined with reinforcement learning, have enabled models to learn from self-play and achieve human parity or even surpass human performance in complex tasks such as the game of Go (Silver et al., 2016, 2017). This naturally raises a question: is it viable to leverage the strengths of MCTS alongside LLMs to inaugurate a novel paradigm of self-improving? More precisely, could the assimilation of MCTS empower LLMs to more effectively explore better responses, guided by strategic signals, and subsequently optimize these responses to enhance overall performance?
@@
To answer this question, we begin with a systematic examination of AlphaGo, identifying three critical aspects for its success: (i) The large volume of expert and self-play data; imitation learning on expert data enables it to simulate human-like strategies, and the reinforcement learning on self-play data fosters the emergence of novel tactics that surpass human capabilities (Clark & Storkey, 2015). (ii) The use of tree search, which facilitates the exploration of potential moves through statistical sampling of the large search space. This approach allows AlphaGo to effectively identify and simulate the most promising strategies, thereby making highly informed decisions in the complex and vast decision space (Silver et al., 2016). (iii) Accurate and unambiguous environment feedback; the direct and accurate feedback (win or loss) provided by the game of Go offers a clear and unequivocal learning signal (Silver et al., 2017). The integration of MCTS with LLMs for self-improvement has several challenges: (i) Limited Data: High-quality annotated data for LLMs is generally scarce. Furthermore, how to construct of synthetic data for LLMs training, similar to AlphaGo’s self-play data, remains unclear. (ii) Search Efficiency: The vast number of potential token combinations in natural language tasks results in an exponentially large search space, posing a significant challenge to the efficiency of MCTS (Ramamurthy et al., 2022). (iii) Imperfect Feedback: In contrast to the clear win/loss feedback in Go, feedback in natural language tasks is often subjective and nuanced, without a straightforward measure of success.
@@
In this paper, we introduce AlphaLLM, an imagination-searching-criticizing framework designed for the self-improvement of LLMs . AlphaLLM consists of three key components, as illustrated in Figure 1. First, an imagination component is designed to synthesize prompts, alleviating the issues of data scarcity. Second, we propose $\eta$Mcts tailored for efficient searching in language tasks. Particularly, it has been show that planning at multiple levels of temporal abstraction is critical for RL problems with a long horizon and large action space (Sutton et al., 1999b; Peng et al., 2017; Luketina et al., 2019). As such, we propose formulating the text generation process as options over a Markov Decision Process (MDP) problem, where each option represents the generation of a collection of tokens for a specific subtask, similar to the concept of chains in chain-of-thought prompting. This formulation improves search efficiency by substantially reducing the search depth. Additionally, we propose the use of state fusion and adaptive branching factors to further enhance search efficiency by balancing the trade-off between search width and depth. Lastly, since accurate feedback is crucial to the success of MCTS, we introduce a trio of critic models to guide $\eta$Mcts, including a value function for estimating future rewards, a process reward model for assessing node correctness, and an outcome reward model for evaluating the overall trajectory. For complex tasks with which LLMs struggle assessing such as arithmetic computation and code execution, to ensure the accuracy of feedback, we augment the critics with the capacity to make dynamic decisions on which tools to use, when to use them, and how to use them effectively. After $\eta$Mcts stage, we collect the trajectory with the largest reward from the critic models as the training examples to improve LLMs.
The experimental results on mathematical reasoning tasks demonstrate that AlphaLLM can efficiently search for better responses and use them to improve LLMs’ performance, forming an effective self-improving loop. Notably, based on LLaMA-2 70b, AlphaLLM can improve its performance from 57.8 to 92.0 on GSM8K and from 20.7 to 51.0 on MATH, performing comparably to GPT-4. In summary, our contributions are threefold:
  • We examine the inherent challenges in harnessing AlphaGo’s self-learning algorithms for LLMs, which are data scarcity, the complexity of search spaces, and the nuanced nature of feedback.
  • We introduce AlphaLLM, an imagination-searching-criticizing framework that integrates MCTS with LLMs, enabling them to self-improve without the need for additional annotations
  • Experiments on mathematical reasoning problems show that, by employing AlphaLLM, we can significantly enhance the performance of LLaMA-2 70B, elevating it to levels comparable with GPT-4 on the GSM8K and MATH datasets when $\eta$Mcts decoding is utilized.
@@ 2 Related Work
Effective search strategy has been shown crucial for tasks that involve complex reasoning and planning, such as go (Silver et al., 2016) and math reasoning (Cobbe et al., 2021; Hendrycks et al., 2021). For math reasoning tasks, various search methods have been studied. One direction of research (Zhu et al., 2024; Xie et al., 2024) designed beam search with dynamic pruning, where beam items of low quality are pruned. Another line of work (Yao et al., 2024; Long, 2023; Besta et al., 2024; Hao et al., 2023; Feng et al., 2023) maintains a tree or a graph that represents the current progress of solving the input question where potential branches are iteratively expanded. Both our approach and Feng et al. (2023) are based on the MCTS algorithm, while one main difference is how to define a search step: Feng et al. (2023) fix a search step to be either a token or a sentence, while our approach is more flexible on deciding steps. More importantly, we also study how to leverage MCTS for effective self-improve. We also design the MCTS process more carefully, such as we merge multiple critique signals to effectively guide the search process. As the result, our approach achieves much better performances than Feng et al. (2023).
@@
Being a key to the success of scalable oversight (Bowman et al., 2022), self-improving for LLM aims to align the LLM to human preference and values mainly using the supervision from the knowledge inside the LLM. One crucial part of self-improving is how to obtain reliable signal of critique to distinguish between good responses from the LLM and bad ones. Initial work (Bai et al., 2022; Wang et al., 2022) first asks the LLM to generate input queries of diverse tasks and the corresponding outputs. They then rely on hand-crafted heuristic rules to filter out redundant or low-quality data pairs (e.g. the query is too long or too short). Since it is non-trivial to compose effective heuristic rule, later work (Sun et al., 2023; Li et al., 2023; Guo et al., 2024) proposes a few general principles or judging criteria and ask the LLM itself to evaluate the quality its responses based on these guidance. They hope that the LLM can automatically designate these principles into each data point to better guide data filtering. However, this requires the LLM to have strong abilities to apply these principles for each specific case and make correct judgements. Different from previous work, we propose to leverage the supervision from MCTS for LLM self-improvement: taking the outputs of MCTS to continue train the LLM. This is because the outputs from MCTS are usually in much better quality then standard nucleus sampling, and the large gap ensure that the LLM can self improve.
Another line of research explores cheaply available knowledge. Some (Saunders et al., 2022; Wang et al., 2023b) collects large-scale critique data from question-and-answer websites (e.g., stack exchange) for continue pretraining, while others (Gou et al., 2023a) utilize external tools to provide more fine-grained guidance. The goal of both directions is to enhance critique ability of the LLM for self-improving. Our approach based on MCTS is intuitively orthogonal to this line of research.
@@ 3.1 Problem Formulation
In this paper, we consider a LLM characterized by probability $p_{\theta}$ and denoted as policy $\pi_{\theta}$. It takes a sequence ${\bm{x}}=[x_{1},\cdots,x_{n}]$ as input, which is typically referred as prompt, to generate the response ${\bm{y}}=[y_{1},\cdots,y_{m}]$. The response ${\bm{y}}$ can be viewed as samples from the conditional probability distribution $p_{\theta}(\cdot|{\bm{x}})$. In the context of LLMs, each $x_{i}$ and $y_{i}$ represents a token from a pre-defined vocabulary. The policy $\pi_{\theta}$ operates in an autoregressive manner, where each token is generated sequentially, relying solely on the context provided by the previously generated tokens. The policy therefore constitutes a Markov process in which the conditional probability distribution $p_{\theta}({\bm{y}}|{\bm{x}})$ can be decomposed and expressed with the chain rule:
$p_{\theta}({\bm{y}}|{\bm{x}})=\prod_{i=1}^{m}p_{\theta}(y_{i}|{\bm{x}},{\bm{y}}_{<i})$
With this property, the text generation task can be formulated as an Markov Decision Process (MDP) problem consisting of $({\mathcal{S}},{\mathcal{A}},T,R,\gamma)$  in which:
  • State ${\bm{s}}_{t}\in{\mathcal{S}}$: Represents the context information of current trajectory, i.e., current status of the generation process, e.g., a partial response to a prompt. The initial state $s_{0}$ corresponds to the original prompt.
  • Action $a_{t}\in{\mathcal{A}}$: Denotes a single action or sampled token from the vocabulary, leading to a transition to a new state ${\bm{s}}_{t+1}$, by concatenating ${\bm{s}}_{t}$ and $a_{t}$.
  • Reward $r_{t}=R({\bm{s}}_{t},a_{t})$: Manifest the evaluation of the generation to the prompt, reflecting the desirability or preferences of each state-action pair, such as whether the actions follow instructions in the prompt.
$\gamma$ denotes the discount factor, while $T$ here signifies the transition probability function. We omit its detailed description as in text generation environment the transition is deterministic.
This MDP framework sets the stage for applying Reinforcement Learning (RL) methods to optimize the policy $\pi_{\bm{\theta}}$ aiming to maximize the expected cumulative reward $R$. Base on these setups, we describe the self-improving problem. Given a LLM $\pi_{\bm{\theta}}$ and an initial dataset ${\mathcal{D}}^{0}$, which consists of $N$ expert-generated prompt-response pairs $\{({\bm{x}}_{i}^{0},{\bm{y}}_{i}^{0})\mid i\in[N]\}$, the goal of self-improving is to iteratively refine $\pi_{\theta}$ to maximize the reward. The refinement process includes learning from synthesized prompts and corresponding responses. These responses are obtained using an advanced search algorithm that navigates the space of possible responses to maximize the expected reward. The detailed process is described in Algorithm 1. The primary challenges in forming an effective self-improving loop lie in synthesizing suitable prompts, efficiently searching over a vast action space, and obtaining precise feedback, which will be discussed in §4.
Output $\theta^{k}$
for $k\leftarrow 1,\dots,K$ do
      Construct dataset ${\mathcal{D}}^{k}=\{({\bm{x}}^{k},\hat{{\bm{y}}}^{k})\}$
end for
Algorithm 1 LLM self-improving loop
@@ 3.2 Monte Carlo Tree Search
MCTS is a sampling-based search algorithm for policy optimization in decision-making problems. It would iteratively build a search tree, by repeating four phases: selection, expansion, evaluation, and backpropagation. In the selection phase, it would recursively select the children from the root node by Upper Confidence Bound (UCB) bandit  Auer et al. (2002), which is
$UCB(i)=w_{i}+C*\sqrt{2*\ln{\frac{N_{i}}{n_{i}}}}$ (1)
where $n_{i}$ and $N_{i}$ are the visit counts for the node $i$ and its parent respectively, $C$ represents a hyperparameter balancing exploration and exploitation, and the $w_{i}$ is the average value of all descendant nodes of $i$. Following selection, the tree undergoes expansion according to the defined policy in the expansion phase. Then in the evaluation phase, the value of the newly expanded node is estimated, by sampling or model-based methods. Finally, in the backpropagation phase, the estimated value is backpropagated to all ancestor nodes of the newly expanded node.

### v2

@@ Abstract
Despite the impressive capabilities of Large Language Models (LLMs) on various tasks, they still struggle with scenarios that involves complex reasoning and planning. Self-correction and self-learning emerge as viable solutions, employing strategies that allow LLMs to refine their outputs and learn from self-assessed rewards. Yet, the efficacy of LLMs in self-refining its response, particularly in complex reasoning and planning task, remains dubious. In this paper, we introduce AlphaLLM for the self-improvements of LLMs, which integrates Monte Carlo Tree Search (MCTS) with LLMs to establish a self-improving loop, thereby enhancing the capabilities of LLMs without additional annotations. Drawing inspiration from the success of AlphaGo, AlphaLLM addresses the unique challenges of combining MCTS with LLM for self-improvement, including data scarcity, the vastness search spaces of language tasks, and the subjective nature of feedback in language tasks. AlphaLLM is comprised of prompt synthesis component, an efficient MCTS approach tailored for language tasks, and a trio of critic models for precise feedback. Our experimental results in mathematical reasoning tasks demonstrate that AlphaLLM significantly enhances the performance of LLMs without additional annotations, showing the potential for self-improvement in LLMs. The code is available at https://github.com/YeTianJHU/AlphaLLM.
@@ 1 Introduction
LLMs, trained on trillions of tokens with billions of parameters have shown unparalleled capabilities in a wide range of natural language processing tasks (Touvron et al., 2023b; Team et al., 2023; OpenAI, 2023). Nevertheless, they continue to face challenges in scenarios requiring complex reasoning and strategic planning  (Valmeekam et al., 2022; Stechly et al., 2024). While advanced prompting approaches such as Chain, Tree, Graph-of-Thought (Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Ding et al., 2023), it remains essential to fine-tune LLMs using a substantial volume of high-quality, supervised data to fundamentally improve the model performance (Nye et al., 2021; Lewkowycz et al., 2022; Chung et al., 2022). This methodology is inherently limited by the scope and quality of data that humans can provide.
@@
Considering these challenges, the concept of self-correction and self-learning have been proposed as promising solutions (Madaan et al., 2024; Saunders et al., 2022; Chen et al., 2024). Within these framework, LLMs typically operate by employing two main strategies: 1) they continuously refine their responses based on the feedback of their past responses, and 2) they extensively sample responses then learn from preferences judged by itself as reward models with PPO or DPO (Yuan et al., 2024a, b; Chen et al., 2024). However, it remains a matter of ongoing research whether LLMs can effectively critique their own outputs to either enhance response quality or apply a scalar reward to indicate the quality of responses, especially in contexts demanding intricate planning and reasoning (Valmeekam et al., 2022; Stechly et al., 2024; Huang et al., 2023; Hong et al., 2023). On the other hand, advanced search algorithms such as MCTS, combined with reinforcement learning, have enabled models to learn from self-play and achieve human parity or even surpass human performance in complex tasks such as the game of Go (Silver et al., 2016, 2017). This naturally raises a question: is it viable to leverage the strengths of MCTS alongside LLMs to inaugurate a novel paradigm of self-improving? More precisely, could the assimilation of MCTS empower LLMs to more effectively explore better responses, guided by strategic signals, and subsequently optimize these responses to enhance overall performance?
@@
To answer this question, we begin with a systematic examination of AlphaGo, identifying three critical aspects for its success: (i) The large volume of data, including self-play data. (ii) The use of tree search, which facilitates the exploration of potential moves through statistical sampling of the large search space. (iii) Accurate and unambiguous environment feedback; the direct and accurate feedback (win or loss) provided by the game of Go offers a clear and unequivocal learning signal (Silver et al., 2017). The integration of MCTS with LLMs for self-improvement has several challenges: (i) Limited Data: High-quality annotated data for LLMs is generally scarce. Furthermore, how to construct of synthetic data for LLMs training, similar to AlphaGo’s self-play data, remains unclear. (ii) Search Efficiency: The vast number of potential token combinations in natural language tasks results in an exponentially large search space, posing a significant challenge to the efficiency of MCTS (Ramamurthy et al., 2022). (iii) Imperfect Feedback: In contrast to the clear win/loss feedback in Go, feedback in natural language tasks is often subjective and nuanced, without a straightforward measure of success.
@@
In this paper, we introduce AlphaLLM, an imagination-searching-criticizing framework designed for the self-improvement of LLMs . AlphaLLM consists of three key components, as illustrated in Figure 1. First, an imagination component is designed to synthesize prompts, alleviating the issues of data scarcity. Second, we propose $\eta$Mcts tailored for efficient searching in language tasks. Particularly, it has been show that planning at multiple levels of temporal abstraction is critical for RL problems with a long horizon and large action space (Sutton et al., 1999b; Peng et al., 2017; Luketina et al., 2019). As such, we propose formulating the text generation process as options over a Markov Decision Process (MDP) problem, where each option represents the generation of a collection of tokens for a specific subtask, similar to the concept of chains in chain-of-thought prompting. This formulation improves search efficiency by substantially reducing the search depth. Additionally, we propose the use of state merge and adaptive branching factors to further enhance search efficiency by balancing the trade-off between search width and depth. Lastly, since accurate feedback is crucial to the success of MCTS, we introduce a trio of critic models to guide $\eta$Mcts, including a value function for estimating expected rewards, a process reward model for assessing node correctness, and an outcome reward model for evaluating the overall trajectory. For complex tasks with which LLMs struggle assessing such as arithmetic computation and code execution, to ensure the accuracy of feedback, we augment the critics with the capacity to make dynamic decisions on which tools to use, when to use them, and how to use them effectively. After $\eta$Mcts stage, we collect the trajectory with the largest reward from the critic models as the training examples to improve LLMs.
The experimental results on mathematical reasoning tasks demonstrate that AlphaLLM can efficiently search for better responses and use them to improve LLMs’ performance, forming an effective self-improving loop. Notably, based on Llama-2-70b and WizardMath-70B-V1.0, AlphaLLM can improve its performance from 57.8 to 92.0 on GSM8K and from 20.7 to 51.0 on MATH, performing comparably to GPT-4.
@@ 2 Related Work
Effective search strategy has been shown crucial for tasks that involve complex reasoning and planning, such as go (Silver et al., 2016) and math reasoning (Cobbe et al., 2021; Hendrycks et al., 2021). For math reasoning tasks, various search methods have been studied. One direction of research (Zhu et al., 2024; Xie et al., 2024) designed beam search with dynamic pruning, where beam items of low quality are pruned. Another line of work (Yao et al., 2024; Long, 2023; Besta et al., 2024; Hao et al., 2023; Feng et al., 2023) maintains a tree or a graph that represents the current progress of solving the input question where potential branches are iteratively expanded. Both our approach and Feng et al. (2023) are based on the MCTS algorithm, while one main difference is how to define a search step: Feng et al. (2023) fix a search step to be either a token or a sentence, while our approach is more flexible on deciding steps. We have also carefully designed the MCTS process, incorporating multiple critique signals to guide the search more effectively and introducing adaptive search parameters for improved state exploration. As the result, our approach achieves much better performances.
@@
Being a key to the success of scalable oversight (Bowman et al., 2022), self-improving for LLM aims to align the LLM to human preference and values mainly using the supervision from the knowledge inside the LLM (Zelikman et al., 2022, 2024). One crucial part of self-improving is how to obtain reliable signal of critique to distinguish between good responses from the LLM and bad ones. Initial work (Bai et al., 2022; Wang et al., 2022) first asks the LLM to generate input queries of diverse tasks and the corresponding outputs. They then rely on hand-crafted heuristic rules to filter out redundant or low-quality data pairs (e.g. the query is too long or too short). Since it is non-trivial to compose effective heuristic rule, later work (Sun et al., 2023; Li et al., 2023; Guo et al., 2024) proposes a few general principles or judging criteria and ask the LLM itself to evaluate the quality its responses based on these guidance, hoping that LLMs can automatically designate these principles into each data point to better guide data filtering. However, this requires LLMs to have strong abilities to apply these principles for each specific case and make correct judgements. Different from previous work, we propose to leverage the supervision from MCTS for LLM self-improvement: taking the outputs of MCTS to continue train the LLM. This is because the outputs from MCTS are usually in much better quality then standard nucleus sampling, and the large gap ensure that the LLM can self improve.
@@ 3.1 Problem Formulation
In this paper, we consider a LLM characterized by probability $p_{\theta}$ and denoted as policy $\pi_{\theta}$. It takes a sequence ${\bm{x}}=[x_{1},\cdots,x_{n}]$ as input, which is typically referred as prompt, to generate the response ${\bm{y}}=[y_{1},\cdots,y_{m}]$. In the context of LLMs, each $x_{i}$ and $y_{i}$ represents a token from a pre-defined vocabulary. The policy $\pi_{\theta}$ operates in an autoregressive manner, where each token is generated sequentially, relying solely on the context provided by the previously generated tokens. The policy therefore constitutes a Markov process in which the conditional probability distribution $p_{\theta}({\bm{y}}|{\bm{x}})$ can be decomposed and expressed with the chain rule as $p_{\theta}({\bm{y}}|{\bm{x}})=\prod_{i=1}^{m}p_{\theta}(y_{i}|{\bm{x}},{\bm{y}}_{<i})$.
With this property, the text generation task can be formulated as an Markov Decision Process (MDP) problem consisting of $({\mathcal{S}},{\mathcal{A}},T,R,\gamma)$  in which, ${\bm{s}}_{t}\in{\mathcal{S}}$ represents the context information of current trajectory, i.e., current status of the generation process, e.g., a partial response to a prompt; $a_{t}\in{\mathcal{A}}$ denotes a single action or sampled token from the vocabulary, leading to a transition to a new state ${\bm{s}}_{t+1}$, by concatenating ${\bm{s}}_{t}$ and $a_{t}$; $r_{t}=R({\bm{s}}_{t},a_{t})$ manifest the evaluation of the generation to the prompt, reflecting the desirability or preferences of each state-action pair.
This MDP framework sets the stage for applying Reinforcement Learning (RL) methods to optimize the policy $\pi_{\bm{\theta}}$ aiming to maximize the expected cumulative reward $R$. Base on these setups, we describe the self-improving problem. Given a LLM $\pi_{\bm{\theta}}$ and an initial dataset ${\mathcal{D}}^{0}$, which consists of $N$ expert-generated prompt-response pairs $\{({\bm{x}}_{i}^{0},{\bm{y}}_{i}^{0})\mid i\in[N]\}$, the goal of self-improving is to iteratively refine $\pi_{\theta}$ to maximize the reward. The refinement process includes learning from synthesized prompts and corresponding responses. These responses are obtained using an advanced search algorithm that navigates the space of possible responses to maximize the expected reward. The detailed process is described in Algorithm 1 in Appendix. The primary challenges in forming an effective self-improving loop lie in synthesizing suitable prompts, efficiently searching over a vast action space, and obtaining precise feedback, which will be discussed in §4.
@@ 3.2 Monte Carlo Tree Search
MCTS is a sampling-based search algorithm for policy optimization in decision-making problems. It would iteratively build a search tree, by repeating four phases: selection, expansion, evaluation, and backpropagation. In the selection phase, it would recursively select the children from the root node by Upper Confidence Bound (UCB)  (Auer et al., 2002), $UCB(i)=w_{i}+C*\sqrt{2*\ln{\frac{N_{i}}{n_{i}}}}$, where $n_{i}$ and $N_{i}$ are the visit counts for the node $i$ and its parent respectively, $C$ represents a hyperparameter balancing exploration and exploitation, and the $w_{i}$ is the average value of all descendant nodes of $i$.